# A Cognitive Evaluation Benchmark of Image Reasoning and Description for Large Vision-Language Models

### v3

@@ Appendix A Image Annotation Instruction
You are going to see some pictures. Each picture tells a story and requires different kinds of reasoning to fully understand the story behind it. You will be first asked to identify the entities and reasoning processes in the picture. Then, you will need to describe the story of the picture based on your identified entities and reasoning processes. First, you will be asked to identify the entities in the picture. The annotation format is [A, B, C], where A, B, C are entities. [Entities]: Please list all entities appearing in the picture, including people, animals, objects etc. You are encouraged to list as many entities as possible. Note that these entities need to be in your picture description afterwards. For entities that are difficult to recognize, please do not list them here or describe them. Then, you will be asked to identify different reasoning processes in the picture. The annotation format should follows the structure [A1 + A2 -> B], where A1 and A2 are premises and B is the conclusion. Note that if you write a conclusion, there must be at least one premise. Do not write a conclusion only, like [B]. Please write one conclusion at a time, do not write a reasoning process like [A1->B->C], which should be split into two. Each picture does not necessarily requires all the reasoning. Please write None, if a picture does not involve a specific kind of reasoning or it is not important in the picture. [Special Time Reasoning] Please write your reasoning processes about the special time of the story in the picture, e.g. festivals, seasons etc. The special time is usually relevant to the story of the picture. For instance, if it is daytime in a picture, it is easily recognized and requires no reasoning and there is nothing special, you can write None. However, if there is a lamp on or a clock indicating a specific time, you can write down your reasoning about time. [Location Reasoning] Please write your reasoning processes about the location of the story in the picture, e.g. near the school. [Character Reasoning]: Please write your reasoning processes about the character of the subjects in the picture, e.g. teacher, doctor etc. [Character Relationship Reasoning]: Please write your reasoning processes about relationship between characters in the picture, e.g. friendship. [Event Reasoning]: Please write your reasoning processes about events happened in the current moment and previous moments in the picture based on clues in the picture. Note that you only need to annotate those high-level events and you can ignore those low-level events. For instance, “the woman is looking at the man” is a low level event and you can ignore its reasoning process. Differently, the reasoning process [A mother is busy cooking. + A boy is fetching cookies behind the mom. + A girl is shushing the boy. -> The boy is stealing cookies.] is a reasoning about high-level event stealing and you should write it down. [Event Relationship Reasoning]: Please write your reasoning processes about relationship between different events in the picture. These events are usually linked through causal and temporal relations. Note that events in this part do not necessarily appears in the [Event Reasoning] part as some events here are low-level events. [Next Moment Event Reasoning]: Please write your reasoning processes about the event that will happen in the next moment. Note that you only need to write down events that have a very high probability of happening, instead of guessing what might happen next. [Mental State Reasoning]: Please write your reasoning processes about mental state of subjects in the picture, e.g. daydreaming, happy etc. You need to reason as best you can about the mental states of all the subjects in the picture, unless they are not showing obvious emotions. Finally, you will be asked to describe the picture in as much detail as you can. [Description]: Please describe all you see in the picture in a paragraph based on the entities and reasoning processes you identified above, ensuring that all of them are included in your description. Each picture has a story behind it and you need to tell that story through your description.
@@ Appendix B Prompt of CoR-based GPT-assisted Question Generation
Figure 7 shows an example prompt of CoR-based GPT-assisted Question Generation for GPT-4. This prompt is used to generate questions based on [Event Reasoning] CoRs. Prompts for other CoR types are similar to this one.
@@
Figure 7: An example prompt of CoR-based GPT-assisted Question Generation for GPT-4 to generate questions based on [Event Reasoning].
Appendix C Prompt of GPT-based Description Cognition Evaluation
Figure 8 and Figure 9 shows prompts of ChatGPT or GPT-4 for cognition evaluation of Description task.
For CoR types other than [Event Relationship Reasoning], we task GPT-4 with determining whether the conclusion in each CoR is mentioned in the description. The prompt is shown in Figure 8. For [Event Relationship Reasoning], we task GPT-4 with determining whether the causal relationship between events (i.e. the whole CoR), as annotated, is present in the description. The prompt is shown in Figure 9.
1) Instead of reasoning about whether each <KEY POINT> is possibly correct based on the <DESCRIPTION>, you only need to determine whether the <DESCRIPTION> mentions the semantics in the <KEY POINT>. 2) Do not overlook the semantics in the <DESCRIPTION> that are semantically equivalent to the <KEY POINT> but expressed in different ways. For instance, if the <DESCRIPTION> mentions "The woman is playing with her son…", we can tell it successfully includes semantics in the <KEY POINT> "The woman is the mother of the boy." 3) If several possible scenarios are listed using ’or’ at a <KEY POINT>, you only need to determine whether one of these scenarios is mentioned in the <DESCRIPTION>. Assign a score of 0 or 1 to each <KEY POINT>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by models.} <KEY POINT>: 1. {Annotated key point 1.} 2. {Annotated key point 2.} N. {Annotated key point N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <KEY POINT>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 8: Evaluation prompt for GPT-4 of Reasoning types other than [Event Relationship Reasoning].
Given a <DESCRIPTION> and some <EVENT RELATIONSHIP>s, please tell me whether this <DESCRIPTION> clearly depicts the cause-and-effect relationships between events. The format of a <EVENT RELATIONSHIP> follows the structure "A1 + A2 + … + An -> B", where A1, A2, …, An and B are events. Events A1, A2, …, An are the causes of event B, and event B is the result caused by events A1, A2, …, An. The criteria for judgment lie in whether the <DESCRIPTION> mentions these events and clearly depicts the causal relationships between them. Assign a score of 0 or 1 to each <EVENT RELATIONSHIP>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by models.} <EVENT RELATIONSHIP>: 1. {Annotated event relationship 1.} 2. {Annotated event relationship 2.} N. {Annotated event relationship N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <EVENT RELATIONSHIP>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 9: Evaluation prompt for GPT-4 of [Event Relationship Reasoning].
Appendix D Introduction to Selected LVLMs
Table 6: LVLMs evaluated in this paper.
@@ Appendix D Introduction to Selected LVLMs
InstructBLIP-7B EVA-CLIP ViT-g Vicuna-7B
Qwen-VL-Chat OpenCLIP ViT-G Qwen-7B
LLaVA-v1.5-7B CLIP ViT-L Vicuna-7B
LLaVA-v1.5-13B CLIP ViT-L Vicuna-13B
mPLUG-Owl-2 CLIP ViT-L LLaMA-7B
ShareGPT4V-7B CLIP ViT-L Vicuna-7B
ShareGPT4V-13B CLIP ViT-L Vicuna-13B
@@
  • InstructBLIP [20] builds upon BLIP-2 [34]. It consists of an image encoder, a LLM, and a Query Transformer (Q-Former). During instruction tuning, only the Q-Former is updated. We use “blip2-vicuna-instruct” for testing.
@@
  • Qwen-VL-Chat [21] is a instruction-tuned VL chatbot based on Qwen-VL. Its training process consists of two stages of pre-training and a final stage of instruction fine-tuning. As for architecture, it consists of a vision encoder, a LLM, and position-aware vision-language adapter. We test “Qwen-VL-Chat”.
@@
  • LLaVA v1.5 [22] is an upgraded version of LLaVA [5]. LLaVA connects a vision encoder and LLM for general-purpose visual and language understanding. It is instruction-tuned on the language-image instruction-following data generated by language-only GPT-4 [1]. By using CLIP-ViT-L-336px with an MLP projection and adding academic-task-oriented VQA data with simple response formatting prompts, LLaVA v1.5 achieves better performance. “llava-v1.5-7b” and “llava-v1.5-13b” are tested.
  • mPLUG-Owl-2 [6] mainly comprises a fundamental vision encoder, a visual abstractor, and a language decoder. It also adopts a two-stage training strategy, comprising pre-training and visual instruction tuning. We test “mplug-owl2-llama2-7b”.
@@
  • ShareGPT4V [23] follows the design of LLaVA v1.5. They incorporate a large-scale resource featuring highly descriptive captions into both the pre-training and supervised fine-tuning phases of ShareGPT4V model. We test“ShareGPT4V-7B” and “ShareGPT4V-13B”.
  • GPT-4V [1] is one of the most powerful LVLMs in the world developed by OpenAI. The version of “gpt-4-vision-preview” is tested.
@@
  • GPT-4o [1] is currently OpenAI’s latest and most powerful multimodal model. It is a single model trained end-to-end across text, vision, and audio. The version of “gpt-4o” is tested.
Table 6 shows an overview of the design of different LVLMs.
Appendix E Evaluation Results of Description Task based on Traditional Image Captioning Metrics
Table 7: Model performance on Description task evaluated by traditional image captioning evaluation metrics.
@@ Appendix E Evaluation Results of Description Task based on Traditional Image Captioning Metrics
InstructBLIP-7B 0.130 0.043 0.255 0.127 0.063 0.033
Qwen-VL-Chat 0.128 0.036 0.232 0.121 0.059 0.031
LLaVA-V1.5-7B 0.146 0.054 0.309 0.158 0.076 0.037
LLaVA-V1.5-13B 0.146 0.051 0.312 0.160 0.076 0.037
mPLUG-Owl-2 0.132 0.035 0.260 0.126 0.057 0.027
ShareGPT4V-7B 0.162 0.017 0.259 0.120 0.050 0.024
ShareGPT4V-13B 0.165 0.024 0.278 0.129 0.055 0.026
GPT-4V 0.208 0.001 0.234 0.115 0.052 0.025
GPT-4o 0.210 0.008 0.254 0.128 0.061 0.030

### v4

@@ Appendix A Image Annotation Instruction
You are going to see some pictures. Each picture tells a story and requires different kinds of reasoning to fully understand it. You will be first asked to identify the entities and reasoning processes in the picture. Then, you will need to describe the story of the picture based on your identified entities and reasoning processes. First, you will be asked to identify the entities in the picture. The annotation format is [A, B, C], where A, B, C are entities. [Entities]: Please list the entities appearing in the picture, including people, animals, objects etc. You are encouraged to list as many entities as possible. Note that these entities need to be in your picture description afterwards. For entities that are difficult to recognize, please do not list them here or describe them. Then, you will be asked to identify different reasoning processes in the picture. The annotation format should follows the structure [A1 + A2 -> B], where A1 and A2 are premises and B is the conclusion. Note that if you write a conclusion, there must be at least one premise. Do not write a conclusion only, like [B]. Please write one conclusion at a time, and do not write a reasoning process like [A1 -> B -> C], which should be split into two. Each picture does not necessarily requires all kinds of reasoning. Please write None, if a picture does not involve a specific kind of reasoning or it is not important in the picture. [Special Time Reasoning] Please write your reasoning processes about the special time of the story in the picture, e.g. festivals, seasons etc. The special time is usually relevant to the story of the picture. For instance, if it is daytime in a picture, it is easily recognized, requires no reasoning and there is nothing special, you can write None. However, if there is a lamp on or a clock indicating a specific time, you can write down your reasoning about the time. [Location Reasoning] Please write your reasoning processes about the location of the story in the picture, e.g. near the school. [Character Reasoning]: Please write your reasoning processes about the characters of the subjects in the picture, e.g. a teacher, a doctor etc. [Character Relationship Reasoning]: Please write your reasoning processes about the relationships between the characters in the picture, e.g. mother-son relationship. [Event Reasoning]: Please write your reasoning processes about the events in the current and previous moments in the picture based on the clues provided. Note that you only need to annotate those high-level events and can ignore the low-level ones. For instance, “the woman is looking at the man” is a low-level event and you can ignore its reasoning process. Differently, the reasoning process [A mother is busy cooking. + A boy is fetching cookies behind the mom. + A girl is shushing the boy. -> The boy is stealing cookies.] is about a high-level event “stealing” and you should write it down. [Event Relationship Reasoning]: Please write your reasoning processes about the relationships between different events in the picture. These events are usually linked through causal and temporal relations. Note that events in this part do not necessarily appears in the [Event Reasoning] part as some events here are low-level events. [Next Moment Event Reasoning]: Please write your reasoning processes about the events that will happen in the next moment. Note that you only need to write down events that have a very high probability of happening, instead of guessing what might happen next. [Mental State Reasoning]: Please write your reasoning processes about the mental states of the subjects in the picture, e.g. daydreaming, happy, etc. You need to reason as best you can about the mental states of all the subjects in the picture, unless they are not showing obvious emotions. Finally, you will be asked to describe the picture in as much detail as you can. [Description]: Please describe all you see in the picture in a paragraph based on the entities and reasoning processes you identified above, ensuring that all of them are included in your description. Each picture has a story behind it and you need to tell that story through your description.
@@ Appendix B Prompt of CoR-based GPT-assisted Question Generation
Figure 7 shows an example prompt of the CoR-based GPT-assisted question generation method for GPT-4. This prompt is used to generate questions based on [Event Reasoning] CoRs. Prompts for other reasoning types are similar to this one.
@@
Figure 7: An example prompt of CoR-based GPT-assisted question generation for GPT-4 to generate questions based on [Event Reasoning] CoRs.
Appendix C Introduction to Selected LVLMs
@@
InstructBLIP-7B EVA-G Vicuna-7B
Qwen-VL-Chat ViT-G/16 Qwen-7B
LLaVA-v1.5-7B CLIP ViT-L/14 Vicuna-v1.5-7B
LLaVA-v1.5-13B CLIP ViT-L/14 Vicuna-v1.5-13B
mPLUG-Owl-2 CLIP ViT-L/14 LLaMA2-7B
ShareGPT4V-7B CLIP ViT-L/14 Vicuna-v1.5-7B
ShareGPT4V-13B CLIP ViT-L/14 Vicuna-v1.5-13B
LLaVA-v1.6-vicuna-7B CLIP ViT-L/14 Vicuna-v1.5-7B
LLaVA-v1.6-vicuna-13B CLIP ViT-L/14 Vicuna-v1.5-13B
LLaVA-v1.6-34B CLIP ViT-L/14 Nous-Hermes-2-Yi-34B
CogVLM-Chat EVA2-CLIP-E Vicuna-v1.5-7B
CogVLM2-Llama3-Chat EVA2-CLIP-E LLaMA3-8B
InternVL2-26B InternViT-6B InternLM2-Chat-20B
Qwen2-VL-7B QwenViT Qwen2-7B
LLaVA-OV-7B-Chat SigLIP-400M Qwen2-7B
@@
  • InstructBLIP Dai et al. (2023) is an extension of BLIP-2 Li et al. (2023b), designed to tackle the challenges of vision-language instruction tuning. It consists of an image encoder, an LLM, and a Q-Former. We use “blip2_vicuna_instruct” + “vicuna7b” for testing.
@@
  • Qwen-VL-Chat Bai et al. (2023) is the instruction-tuned VL chatbot based on Qwen-VL, which consists of a visual encoder, an LLM, and a position-aware vision-language adapter. Its training process consists of two pre-training stages followed by a final instruction fine-tuning stage. We test “Qwen-VL-Chat” in the paper.
@@
  • LLaVA v1.5 (Liu et al., 2023a) is an upgraded version of LLaVA Liu et al. (2023b), who connects a vision encoder and LLM for visual and language understanding. LLaVA is instruction-tuned on the language-image instruction-following data generated by language-only GPT-4. By using CLIP-ViT-L-336px with an MLP projection and adding academic-task-oriented VQA data with simple response formatting prompts, LLaVA v1.5 achieves better performance. “llava-v1.5-7b” and “llava-v1.5-13b” are tested.
  • mPLUG-Owl-2 Ye et al. (2023) leverages modality collaboration to enhance performance across both text and multi-modal tasks. It adopts a modularized network design, with the language decoder serving as a universal interface to manage different modalities. We test “mplug-owl2-llama2-7b” in the paper.
@@
  • ShareGPT4V Chen et al. (2023a) follows the design of LLaVA v1.5. It incorporates a large-scale resource featuring highly descriptive captions into both the pre-training and supervised fine-tuning phases. We test “ShareGPT4V-7B” and “ShareGPT4V-13B”.
  • LLaVA v1.6 Liu et al. (2024) maintains the minimalist design and data efficiency of LLaVA-v1.5. It enhances LLaVA-v1.5 with dynamic high resolution, data mixture, and scaling of the LLM backbone. We test “llava-v1.6-vicuna-7b-hf”, “llava-v1.6-vicuna-13b-hf” and “llava-v1.6-34b-hf” in the paper.
  • CogVLM Wang et al. (2023) comprises a ViT encoder, an MLP adapter, a pretrained LLM, and a visual expert module. Unlike the common shallow alignment method that maps image features to the input space of language model, CogVLM uses a trainable visual expert module in the attention and FFN layers to bridge the gap between the frozen pretrained language model and image encoder. “cogvlm-chat-hf” is tested in the paper.
  • CogVLM2 Hong et al. (2024) inherits the visual expert architecture from CogVLM. Differently, CogVLM2 uses a 2×2 downsampling module to increase input resolution while maintaining efficiency, with LLaMA3-8B as its backbone. The pre-training and post-training data are also improved in terms of diversity and quality. “cogvlm2-llama3-chat-19B” is tested.
  • InternVL2 Chen et al. (2023c) family adopts a progressive alignment training strategy, resulting in the first vision foundation model natively aligned with LLMs, and enabling efficient training of large models with limited resources. It supports multimodal input using a single parameter set and provides multitask output, including images, bounding boxes, and masks. By connecting the LVLM with multiple downstream task decoders, it can be generalized to many VL tasks. We test “InternVL2-26B” in this paper.
  • Qwen2-VL Wang et al. (2024) is built upon the Qwen-VL architecture. It introduces the Naive Dynamic Resolution mechanism, which allows the model to handle arbitrary image resolutions, dynamically adjusting the number of visual tokens. It also incorporates Multimodal Rotary Position Embedding (M-RoPE), enabling the effective fusion of positional information across text, images, and videos. “Qwen2-VL-7B-Instruct” is tested.
  • LLaVA-OneVision-Chat Li et al. (2024) is an enhanced version of LLaVA-OneVision, with improvements achieved through preference alignment for better visual-chat capabilities. LLaVA-OneVision is developed by integrating insights into data, models, and visual representations in the LLaVA v1.6 series. “llava-onevision-qwen2-7b-ov-chat” is tested and we use LLaVA-OV-Chat to refer to LLaVA-OneVision-Chat in this paper.
  • GPT-4V (OpenAI, 2023) is a powerful LVLM developed by OpenAI. The version of “gpt-4-turbo” is tested.
@@
  • GPT-4o (OpenAI, 2023) is currently one of the most powerful multimodal models. It is a single model trained end-to-end across text, vision, and audio. The version of “gpt-4o” is tested.
Table 6 shows an overview of the designs of different LVLMs.
Appendix D Prompts for GPT-based Cognition Evaluation of the Description Task
Figure 8 and Figure 9 show the prompts used for cognition evaluation of the Description task for GPT-4 (or ChatGPT).
1) Instead of reasoning about whether each <KEY POINT> is possibly correct based on the <DESCRIPTION>, you only need to determine whether the <DESCRIPTION> mentions the semantics in the <KEY POINT>. 2) Do not overlook the semantics in the <DESCRIPTION> that are semantically equivalent to the <KEY POINT> but expressed in different ways. For instance, if the <DESCRIPTION> mentions "The woman is playing with her son…", we can tell it successfully includes semantics in the <KEY POINT> "The woman is the mother of the boy." 3) If several possible scenarios are listed using ‘or’ at a <KEY POINT>, you only need to determine whether one of these scenarios is mentioned in the <DESCRIPTION>. Assign a score of 0 or 1 to each <KEY POINT>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by a model.} <KEY POINT>: 1. {Annotated key point 1.} 2. {Annotated key point 2.} N. {Annotated key point N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <KEY POINT>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 8: Cognition evaluation prompt of reasoning types other than [Event Relationship Reasoning].
Given a <DESCRIPTION> and some <EVENT RELATIONSHIP>s, please tell me whether this <DESCRIPTION> clearly depicts the cause-and-effect relationships between events. The format of a <EVENT RELATIONSHIP> follows the structure "A1 + A2 + … + An -> B", where A1, A2, …, An and B are events. Events A1, A2, …, An are the causes of event B, and event B is the result caused by events A1, A2, …, An. The criteria for judgment lie in whether the <DESCRIPTION> mentions these events and clearly depicts the causal relationships between them. Assign a score of 0 or 1 to each <EVENT RELATIONSHIP>, where 0 represents NO and 1 represents YES. <DESCRIPTION>: {Description generated by a model.} <EVENT RELATIONSHIP>: 1. {Annotated event relationship 1.} 2. {Annotated event relationship 2.} N. {Annotated event relationship N.} Please write your answers in “[ ]” with 0 or 1 in the following format (number + square brackets): 1. [1] 2. [0] Your answers to the {N} <EVENT RELATIONSHIP>(s) above: 1. [ ] 2. [ ] … N. [ ]
Figure 9: Cognition evaluation prompt of [Event Relationship Reasoning].
@@ Appendix D Prompts for GPT-based Cognition Evaluation of the Description Task
InstructBLIP-7B 0.130 / 0.183 0.043 / 0.003 0.255 / 0.218 0.127 / 0.104 0.063 / 0.049 0.033 / 0.024
Qwen-VL-Chat 0.130 / 0.196 0.037 / 0.016 0.242 / 0.262 0.124 / 0.138 0.059 / 0.069 0.030 / 0.036
LLaVA-V1.5-7B 0.146 / 0.182 0.054 / 0.020 0.309 / 0.275 0.158 / 0.138 0.076 / 0.065 0.037 / 0.032
LLaVA-V1.5-13B 0.146 / 0.176 0.051 / 0.018 0.312 / 0.274 0.160 / 0.137 0.076 / 0.066 0.037 / 0.034
mPLUG-Owl-2 0.132 / 0.184 0.035 / 0.012 0.260 / 0.235 0.126 / 0.116 0.057 / 0.053 0.027 / 0.025
ShareGPT4V-7B 0.162 / 0.191 0.017 / 0.014 0.259 / 0.222 0.120 / 0.113 0.050 / 0.053 0.024 / 0.027
ShareGPT4V-13B 0.165 / 0.184 0.024 / 0.015 0.278 / 0.254 0.129 / 0.126 0.055 / 0.059 0.026 / 0.030
LLaVA-v1.6-vicuna-7B 0.169 / 0.190 0.026 / 0.000 0.278 / 0.159 0.134 / 0.085 0.060 / 0.041 0.029 / 0.021
LLaVA-v1.6-vicuna-13B 0.172 / 0.195 0.023 / 0.000 0.283 / 0.166 0.137 / 0.090 0.061 / 0.043 0.030 / 0.021
LLaVA-v1.6-34B 0.172 / 0.199 0.026 / 0.000 0.279 / 0.173 0.134 / 0.092 0.060 / 0.044 0.030 / 0.021
CogVLM-Chat 0.157 / 0.174 0.069 / 0.056 0.297 / 0.326 0.151 / 0.154 0.075 / 0.069 0.039 / 0.034
CogVLM2-Llama3-Chat 0.176 / 0.191 0.059 / 0.032 0.356 / 0.257 0.183 / 0.127 0.091 / 0.059 0.049 / 0.029
InternVL2-26B 0.210 / 0.202 0.008 / 0.000 0.250 / 0.166 0.136 / 0.091 0.070 / 0.044 0.037 / 0.023
Qwen2-VL-7B 0.205 / 0.183 0.007 / 0.000 0.222 / 0.154 0.121 / 0.084 0.061 / 0.041 0.032 / 0.021
LLaVA-OV-7B-Chat 0.198 / 0.192 0.009 / 0.000 0.245 / 0.138 0.118 / 0.072 0.051 / 0.033 0.024 / 0.016
GPT-4V 0.189 / 0.191 0.013 / 0.000 0.250 / 0.153 0.113 / 0.071 0.047 / 0.028 0.022 / 0.011
GPT-4o 0.182 / 0.189 0.070 / 0.000 0.332 / 0.145 0.161 / 0.074 0.077 / 0.033 0.039 / 0.016
Table 7: Model performance on the Description task evaluated using traditional image captioning metrics. The results of the evaluation under the Spontaneous Description mode and Directed Reasoning mode are presented before and after the “/” in each table cell.
Appendix E Evaluation of LVLMs on the Description Task Using Traditional Image Captioning Metrics